OpenStudy (anonymous):

when I'm given that f(x) = an algebraic equation and asked to evaluate f(4)...how do I get started?

OpenStudy (anonymous):

substitute all the x's for 4

OpenStudy (anonymous):

for instance f(x)=x^2+4x =4^2+4(x)

OpenStudy (anonymous):

so, if evaluating f(1) where f(x)=-3x^3-3x-6, then 3x^3 is -3 and -3x is -3 and -6 equals -12?

OpenStudy (anonymous):

Thanks...I used your vector to re-blue my old math experience and for that, I thank you warmly, Bowate.